# Does the square-root price impact law belong to the strict universal scalings?: quantitative support by a complete survey of the Tokyo stock exchange market

### v1

@@ Abstract
Universal power laws have been scrutinised in physics and beyond, and a long-standing debate exists in econophysics regarding the strict universality of the nonlinear price impact, commonly referred to as the square-root law (SRL). The SRL posits that the average price impact $I$ follows a power law with respect to transaction volume $Q$, such that $I(Q)\propto Q^{\delta}$ with $\delta\approx 1/2$. Some researchers argue that the exponent $\delta$ should be system-specific, without universality. Conversely, others contend that $\delta$ should be exactly $1/2$ for all stocks across all countries, implying universality. However, resolving this debate requires high-precision measurements of $\delta$ with errors of around $0.1$ across hundreds of stocks, which has been extremely challenging due to the scarcity of large microscopic datasets – those that enable tracking the trading behaviour of all individual accounts. Here we conclusively support the universality hypothesis of the SRL by a complete survey of all trading accounts for all liquid stocks on the Tokyo Stock Exchange (TSE) over eight years. Using this comprehensive microscopic dataset, we show that the exponent $\delta$ is equal to $1/2$ within statistical errors at both the individual stock level and the individual trader level. Additionally, we rejected two prominent models supporting the nonuniversality hypothesis: the Gabaix-Gopikrishnan-Plerou-Stanley and the Farmer-Gerig-Lillo-Waelbroeck models. Our work provides exceptionally high-precision evidence for the universality hypothesis in social science and could prove useful in evaluating the price impact by large investors – an important topic even among practitioners.
@@
Figure 1: Square-root price impact law. When a trader executes a large volume of order (called a metaorder, with size $Q$), they typically split them into a long sequence of child orders (with length $L$). The resulting price impact obeys the power law $\mathcal{I}(\mathcal{Q})\propto c\mathcal{Q}^{\delta}$ with $\delta\approx 1/2$.
@@
Universality is one of the most appealing topics in physics, whereby common macroscopic laws are observed across various materials StanleyB1971 ; GoldenfeldB . For example, universal power laws are observed in critical phenomena, such as $M\propto|T-T_{c}|^{\delta}$, with an order parameter $M$ near the critical point $T_{c}$ (e.g., the magnetisation $M$ and temperature $T$ for a ferromagnetic system). Interestingly, the exponent $\delta$ is independent of the microscopic details of the system. Such a robust exponent is referred to as a universal exponent to emphasise the special significance of universality, in contrast to a nonuniversal exponent that may be a fragile observable, as it could easily vary with perturbations to the system’s microscopic parameters.
@@
Inspired by such success, physicists have attempted to discover universal power laws beyond physics, including in social science. For instance, in finance, some physicists believe the square-root law (SRL) for price impact BouchaudText should be part of such universal scaling laws Toth2011 ; MastromatteoPRL2014 ; MastromatteoPRE2014 ; Bucci2019PRL . However, this claim of universality has not been fully established for lack of large-scale financial microscopic datasets that enable the tracking of all trading accounts individually. Here we quantitatively support the universality hypothesis through a complete survey of all trading accounts on the Tokyo Stock Exchange (TSE), aiming to conclude the long-standing debate in econophysics.
@@
The SRL is an empirical law that describes the nonlinear price response following large buying (selling) by institutional investors. Consider a trader who buys (or sells) $Q$ units of stock. Following this transaction, the market responds with a price movement $\Delta p$. For small $Q$, a linear response is expected, such that $I(Q):=\langle\epsilon\Delta p\mid Q\rangle\approx\lambda Q$, where $I(Q)$ is called the price impact, $\epsilon$ is the order sign defined by $\epsilon=+1$ ($\epsilon=-1$) for buying (selling), and $\langle A\mid B\rangle$ represents the average of $A$ conditioned on $B$. For large $Q$, on the contrary, a nonlinear response is empirically observed (see Fig. 1),
@@
after the nondimensionalisation $\mathcal{Q}:=Q/V_{\rm D}$ and $\mathcal{I}(\mathcal{Q}):=I(Q)/\sigma_{\rm D}$, where $V_{\rm D}$ is the daily transaction volume and $\sigma_{\rm D}$ is the daily volatility. Surprisingly, the exponent $\delta$ is typically claimed to be close to one-half for any asset all over the world,
$\delta\approx\frac{1}{2}$ (2)
@@
and, thus, the nonlinear scaling (1) is referred to as the SRL Toth2011 ; MastromatteoPRL2014 ; MastromatteoPRE2014 ; Bucci2019PRL ; DonierLOB2015 . The SRL closely resembles the universal power law observed in physics and has been an exceptionally appealing topic in econophysics (see the Supplementary information (SI) for a detailed literature review).
@@
However, there has been a long-standing debate regarding the strict universality of the SRL. Some physicists argue that the exponent $\delta$ should depend on the microscopic details of financial markets and, therefore, $\delta$ should be considered a nonuniversal exponent. For instance, the Gabaix-Gopikrishnan-Plerou-Stanley (GGPS) and the Farmer-Gerig-Lillo-Waelbroeck (FGLW) models are two promising microscopic frameworks that support the nonuniversality hypothesis NatureGabaix ; Gabaix2006 ; FGLWQFin2013 . Conversely, other physicists contend that $\delta$ should be exactly one-half for all financial markets in any country, making $\delta$ a universal exponent. For example, the latent order-book model Toth2011 ; MastromatteoPRL2014 ; MastromatteoPRE2014 ; DonierLOB2015 proposes a minimal mechanism to derive $\delta=1/2$, supporting the universality hypothesis. Which hypothesis is ultimately correct?
@@
Despite its evident importance, this debate has not yet been settled for several reasons. First, microscopic financial datasets – those that enable tracking trading behaviour for all individual trading accounts – are essential to test these hypotheses. Unfortunately, such microscopic datasets are extremely rare, making it challenging to rigorously test these hypotheses. Second, while previous studies have investigated the SRL, their datasets were typically proprietary, provided by specific hedge funds. This means that their datasets are neither comprehensive nor randomly sampled, which introduces the possibility of sampling bias, and the resulting sample sizes are inevitably small. A complete survey – studying all trading accounts for all liquid stocks – has been highly desired to address this issue convincingly. Third, solving these issues requires extremely high accuracy in measuring the exponent $\delta$. To test these hypotheses, we need to measure $\delta$ with an accuracy of about $0.1$ for hundreds of stocks. Such a high-precision measurement is difficult to achieve in social sciences for lack of large microscopic datasets.
@@
In this article, we present conclusive evidence supporting the universality of the SRL by a complete survey of the TSE. We analysed a large, high-quality microscopic dataset that records the percise trading behaviour of all trading accounts on the TSE for all stocks over eight years. Our TSE dataset includes virtual server identifiers (IDs) – a unit of trading accounts on the TSE – enabling us to investigate the resulting price impact $\Delta p$ associated with the buy (sell) volume $Q$ at the individual account level. Using this dataset, we precisely measured $\delta$ by controlling errors of the order of $0.07$ and found that $\delta$ was equal to $1/2$ for all liquid stocks (more than two thousand data points) on the TSE within the statistical errors. We also measured $\delta$ for all active traders individually and found it to be equal to $1/2$ within statistical errors as well. Furthermore, we rejected two prominent models that have theoretically supported the nonuniversality hypothesis: the GGPS and the FGLW models. Our result supports the universality hypothesis of the SRL with extraordinary accuracy and suggests that it can be robustly used to estimate transaction costs for institutional investors.
@@
Our dataset was provided by Japan Exchange (JPX) Group, Inc., the platform manager of the TSE market. This dataset records the complete life cycle of all order submissions for all stocks on the TSE from 4th January 2012 to 2nd November 2019. The unique feature of our dataset is that it includes virtual server IDs. Virtual server IDs are not technically equivalent to the membership IDs, because any trader may possess several virtual servers. However, we can define effective trader IDs, referred to as trading desks, by tracking all the virtual server IDs (see Method, the SI, and our previous studies Goshima2019 ; SatoPRL2023 ; SatoPRR2023 ). In this article, we refer to these trading desks as trader IDs for short.
We focus on price impacts during the continuous double auction sessions: from 09:00 to 11:30 and from 12:30 to 15:00 Japan Standard Time. The dataset was divided into two periods: from 4th January 2012 to 18th September 2015 (Dataset 1) and from 24th September 2015 to 2nd November 2019 (Dataset 2). Note that technical updates were made to the trading system (called arrowhead) between the two datasets, and the trader IDs were reallocated. We regard any stock in Dataset 1 as a different stock from Dataset 2, even if the ticker code remains the same. For this definition, we have 942 stocks in Dataset 1 and 1,357 stocks in Dataset 2. Therefore, we can measure $\delta$ for these 2,299 datapoints.
@@
Figure 2: Stock-level price impact. (a) Average price-impact profile for four stocks (red: Toyota Motor Corp., blue: Nippon Telegraph and Telephone (NTT) Corp., green: SoftBank Group Corp., orange: Takeda Pharmaceutical Company Limited). Errorbars represent the standard errors of the means (SEMs). (b) Aggregate scaling plot, visualising the average scaling profiles across all stocks, reveals the validity of the square-root law $\mathcal{I}(\mathcal{Q})\propto c\mathcal{Q}^{1/2}$ for all stocks. (c) Histogram of the exponent $\delta$ across all stocks, exhibiting a sharp peak around $0.5$ with $\langle\delta\rangle=0.489$ with its SEM of $0.0015$ and $\sigma_{\delta}:=\sqrt{\langle(\delta-\langle\delta\rangle)^{2}\rangle}=0.071$. The size of the standard deviation $\sigma_{\delta}$ is at the same level as the finite sample-size effect $\langle\!\langle\sigma_{\delta}\rangle\!\rangle=0.062$, according to our simple price-series model with exact $\delta=1/2$. (d) Histogram of the prefactor $c$ across all stocks, with $\langle c\rangle=0.842$.
@@
We measured the transaction volume $Q$ and corresponding price impact $\Delta p$ for all liquid stocks on the TSE. First, we defined trader IDs for each stock and extracted the market-order volume sequences. Second, we assume that institutional investors typically split their large orders (called metaorders) into small orders (called child orders), as suggested by the literature LMF_PRE2005 ; SatoPRL2023 ; SatoPRR2023 ; Bershova ; GeneralizedLMF . Specifically, if two successive market orders have the same sign (plus for buy and minus for sell), we assume that both orders are child orders belonging to the same metaorder. The total volume $Q$ is then defined as the sum of all child-order volumes. Third, the price impact $\Delta p$ is defined as the price movement caused by such a metaorder. In this article, liquid stocks are defined as stocks where the total number of metaorders exceeds $10^{5}$.
@@
Figure 2(a) presents the typical price-impact profile $I(Q)$ for four liquid stocks. Additionally, we present Fig. 2(b) as the aggregate scaling plot – visualising the averaged impact profile across all stocks as one figure. The aggregate scaling plot was created by two steps (see Methods and the SI): (i) After non-dimensionalising $\mathcal{Q}:=Q/V_{\rm D}$ and $\mathcal{I}(\mathcal{Q}):=I(Q)/\sigma_{\rm D}$ for each stock, we applied fitting $\mathcal{I}(\mathcal{Q})=cQ^{1/2}$ to estimate $c$. (ii) The average of $\mathcal{I}(\mathcal{Q})/c$ across all stocks was plotted, which collapsed onto a single master curve as theoretically expected. These two figures provide an overall picture of the SRL (1) and qualitatively support its validity across all stocks.
We next present the normalised histogram of $\delta$ across all stocks as Fig. 2(c). Based on the SRL (1), we measured $\delta$ and $c$ by the method of relative least squares for all stocks and plotted their histogram. Fig. 2(c) shows a very sharp peak around $1/2$ with the average and the standard deviation,
$\langle\delta\rangle=0.489\pm 0.0015,\>\>\>\sigma_{\delta}:=\sqrt{\langle(\delta-\langle\delta\rangle)^{2}\rangle}=0.071$ (3)
@@
where $\langle A\rangle$ represents the average of $A$ across all stocks. This is the first main result of this article, suggesting that the exact SRL holds for more than two thousand data points, with the dispersion controlled to less than $0.1$. As a reference, we also present the histogram of the prefactor $c$ in Fig. 2(d).
While Fig. 2(c) provides quantitative evidence of the SRL with a precision of $0.071$ for all stocks, one might be interested in understanding the origin of the dispersion $\sigma_{\delta}$ to establish the universality hypothesis of the SRL. For example, if this level of dispersion can be attributed to the finite sample-size effect, we can confidently claim that the universality hypothesis holds within statistical errors, at least on the TSE.
[…]
$\langle\!\langle\delta\rangle\!\rangle=0.489\pm 0.0013,\>\>\>\langle\!\langle\sigma_{\delta}\rangle\!\rangle=0.062\pm 0.0010.$ (4)
@@
Since $\sigma_{\delta}\approx\langle\!\langle\sigma_{\delta}\rangle\!\rangle$, we concluded that the standard deviation in estimating $\delta$ from our dataset is almost attributable to finite sample-size effects, and the universality hypothesis holds for the SRL within statistical errors, at least for the TSE.
@@
Figure 3: Trader-level price impact. (a) Trader-level aggregate scaling plot, visualising the average scaling profile across all traders, reveals the validity of the square-root law for all traders. (b) Histogram of the exponent $\delta^{(i)}$ across all traders, focusing on the range $0\leq\delta^{(i)}\leq 1$. Due to fitting errors, a few outliers were observed outside this region and were excluded. The histogram exhibits a sharp peak around $0.5$, with $\langle\delta^{(i)}\rangle=0.493$ with its SEM of $0.0050$ and $\sigma_{\delta^{(i)}}:=\sqrt{\langle(\delta^{(i)}-\langle\delta^{(i)}\rangle)^{2}\rangle}=0.177$. The size of the standard deviation $\sigma_{\delta^{(i)}}$ is at the same level as the finite sample-size effect $\langle\!\langle\sigma_{\delta^{(i)}}\rangle\!\rangle=0.167$. (c) Histogram of the prefactor $c^{(i)}$ across all traders, with $\langle c^{(i)}\rangle=1.501$.
While we presented evidence at the stock level, is the SRL universal even at the trader level? In other words, by writing the exponent for the $i$-th trader as $\delta^{(i)}$, is $\delta^{(i)}$ statistically indistinguishable from $1/2$ for all traders, despite the diversity in trading strategies among traders? To answer this question, we studied the price impact of all active traders – roughly defined as traders who submitted metaorders not less than $10^{4}$ times in this article (see Method and SI). With this definition, we have 1,293 active traders in total (415 from Dataset 1 and 878 from Dataset 2). We present the aggregate scaling plot at the individual trader level as Fig. 3(a), qualitatively suggesting the validity of the SRL for all traders. As a second main result, we present the normalised trader-level histogram of $\delta^{(i)}$ as Fig. 3(b). The histogram has clear sharp peak with the average $\langle\delta^{(i)}\rangle$ and standard deviation $\sigma_{\delta^{(i)}}:=\sqrt{\langle(\delta^{(i)}-\langle\delta^{(i)}\rangle)^{2}\rangle}$ as
@@
The standard deviation $\sigma_{\delta^{(i)}}$ in our data was at the same level as that evaluated by our numerical model, $\langle\!\langle\delta^{(i)}\rangle\!\rangle=0.521\pm 0.0048$ and $\langle\!\langle\sigma_{\delta^{(i)}}\rangle\!\rangle=0.167\pm 0.0074$. Since $\sigma_{\delta^{(i)}}\approx\langle\!\langle\sigma_{\delta^{(i)}}\rangle\!\rangle$, we concluded that the origin of the standard deviation $\sigma_{\delta^{(i)}}$ is almost attributable to the finite sample-size effect, supporting the universality hypothesis of the SRL within statistical errors even at the trader level.

### v2

# Strict universality of the square-root law in price impact across stocks: a complete survey of the Tokyo stock exchange

@@ Abstract
Universal power laws have been scrutinised in physics and beyond, and a long-standing debate exists in econophysics regarding the strict universality of the nonlinear price impact, commonly referred to as the square-root law (SRL). The SRL posits that the average price impact II follows a power law with respect to transaction volume $Q$, such that $I(Q)\propto Q^{\delta}$ with $\delta\approx 1/2$. Some researchers argue that the exponent $\delta$ should be system-specific, without universality. Conversely, others contend that $\delta$ should be exactly $1/2$ for all stocks across all countries, implying universality. However, resolving this debate requires high-precision measurements of $\delta$ with errors of around $0.1$ across hundreds of stocks, which has been extremely challenging due to the scarcity of large microscopic datasets—those that enable tracking the trading behaviour of all individual accounts. Here we conclusively support the universality hypothesis of the SRL by a complete survey of all trading accounts for all liquid stocks on the Tokyo Stock Exchange (TSE) over eight years. Using this comprehensive microscopic dataset, we show that the exponent $\delta$ is equal to $1/2$ within statistical errors at both the individual stock level and the individual trader level. Additionally, we rejected two prominent models supporting the nonuniversality hypothesis: the Gabaix-Gopikrishnan-Plerou-Stanley and the Farmer-Gerig-Lillo-Waelbroeck models (Nature 2003, QJE 2006, and Quant. Finance 2013). Our work provides exceptionally high-precision evidence for the universality hypothesis in social science and could prove useful in evaluating the price impact by large investors—an important topic even among practitioners.
Introduction.
@@
Figure 1: Square-root price impact law. Practitioners typically split a large volume of market orders (called a metaorder, with size $Q$) into a long sequence of child orders (with length $L$). Price impact $I(Q):=\langle\epsilon\Delta p\mid Q\rangle$ is the average mid-price change $\Delta p$ between the beginning and the end of a buy (sell) metaorder with $\epsilon=+1$ ($\epsilon=-1$) and follows $I(Q)\propto Q^{\delta}$ with $\delta\approx 1/2$.
@@
Universality is one of the central topics in physics, whereby common macroscopic laws are observed across various materials [1, 2]. For example, universal power laws are observed in critical phenomena, such as $M\propto|T-T_{c}|^{\delta}$, with an order parameter $M$ near the critical point $T_{c}$ (e.g., the magnetisation $M$ and temperature $T$ for a ferromagnetic system). Interestingly, the exponent $\delta$ is independent of the microscopic details of the system. Such a robust exponent is referred to as a universal exponent to emphasise the special significance of universality, in contrast to a nonuniversal exponent that may be a fragile observable, as it could easily vary with perturbations to the system’s microscopic parameters.
@@
Inspired by such success, physicists have attempted to discover universal power laws beyond physics, including in social science. For instance, in finance, some physicists believe the square-root law (SRL) for price impact [3] should be part of such universal scaling laws [5, 6, 7, 4, 8]. At the same time, this universality claim has not been fully established for lack of large financial microscopic datasets that enable the tracking of all trading accounts individually. Here we quantitatively support the universality hypothesis through a complete survey of all trading accounts on the Tokyo Stock Exchange (TSE), aiming to conclude the long-standing debate in econophysics.
@@
The SRL is an empirical law that describes the nonlinear price response following large buying (selling) by institutional investors. Consider a trader buying (or selling) $Q$ units of stock by market orders—immediate decisions to buy or sell shares. Following this transaction, the market responds with a price movement $\Delta p$. For small $Q$, a linear response is expected, such that $I(Q):=\langle\epsilon\Delta p\mid Q\rangle\approx\lambda Q$, where $I(Q)$ is called the price impact [3], $\epsilon$ is the market order sign defined by $\epsilon=+1$ ($\epsilon=-1$) for buying (selling), and $\langle A\mid B\rangle$ represents the average of AA conditioned on $B$. For large $Q$, on the contrary, a nonlinear response is empirically observed (see Fig. 1),
@@
after the nondimensionalisation $\mathcal{Q}:=Q/V_{\rm D}$ and $\mathcal{I}(\mathcal{Q}):=I(\mathcal{Q})/\sigma_{\rm D}$, where $V_{\rm D}$ is the daily transaction volume and $\sigma_{\rm D}$ is the daily volatility. Surprisingly, the exponent $\delta$ is typically claimed to be close to one-half for any asset worldwide,
$\delta\approx\frac{1}{2}.$ (2)
@@
Thus, the nonlinear scaling (1) is referred to as the SRL [5, 6, 7, 4, 8]. For example, the SRL predicts buying 1% of a stock’s daily traded shares typically moves the price by about 10% of its daily price range. The SRL closely resembles the universal power law observed in physics and has been an exceptionally appealing topic in econophysics (see Supplemental Material (SM) [9] for a detailed literature review).
@@
However, there has been a long-standing debate regarding the strict universality of the SRL. Some physicists argue that the exponent $\delta$ should depend on the microscopic details of financial markets and, therefore, $\delta$ should be considered a nonuniversal exponent. For instance, the Gabaix-Gopikrishnan-Plerou-Stanley (GGPS) and the Farmer-Gerig-Lillo-Waelbroeck (FGLW) models are two promising microscopic frameworks that support the nonuniversality hypothesis [24, 25, 26]. Conversely, other physicists contend that $\delta$ should be exactly one-half for all financial markets in any country, making $\delta$ a universal exponent. For example, the latent order-book model [5, 6, 7, 8] proposes a minimal mechanism to derive $\delta=1/2$, supporting the universality hypothesis. Which hypothesis is ultimately correct? This is a major point even in financial economics, as hypothesis selection is strongly relevant for validating/rejecting microscopic models to explain the SRL’s mechanisms.
@@
Despite its evident importance, this debate has not yet been settled for several reasons. First, microscopic financial datasets—those that enable tracking trading behaviour for all individual trading accounts—are essential to test these hypotheses. Unfortunately, such microscopic datasets are extremely rare, making it challenging to rigorously test these hypotheses. Second, while previous studies have investigated the SRL, their datasets were typically proprietary, provided by specific hedge funds. This means that their datasets are neither comprehensive nor randomly sampled, which introduces the possibility of sampling bias, and the resulting sample sizes are inevitably small. A complete survey—studying all trading accounts for all liquid stocks—has been highly desired to address this issue convincingly. Third, solving these issues requires extremely high accuracy in measuring the exponent $\delta$. To test these hypotheses, we need to measure $\delta$ with an accuracy of about $0.1$ for hundreds of stocks. Such a high-precision measurement is difficult to achieve in social sciences for lack of large microscopic datasets.
@@
In this article, we present conclusive evidence supporting the universality of the SRL by a complete survey of the TSE. We analysed a large, high-quality microscopic dataset that records the percise trading behaviour of all trading accounts on the TSE for all stocks over eight years. Our TSE dataset includes virtual server identifiers (IDs)—a unit of trading accounts on the TSE—enabling us to investigate the resulting price impact $\Delta p$ associated with the buy (sell) volume $Q$ at the individual account level. Using this dataset, we precisely measured $\delta$ by controlling errors of the order of $0.06$ and found that $\delta$ was equal to $1/2$ for all liquid stocks (more than two thousand data points) within the statistical errors. We also measured $\delta$ for all active traders individually and found it to be equal to $1/2$ within statistical errors as well. Furthermore, we rejected two prominent models that had theoretically supported the nonuniversality hypothesis: the GGPS and the FGLW models. Our result supports the universality hypothesis of the SRL with extraordinary accuracy.
@@
Our dataset was provided by Japan Exchange (JPX) Group, Inc., the platform manager of the TSE market. This dataset records the complete life cycle of all order submissions for all stocks on the TSE over almost eight years. The unique feature of our dataset is that it includes virtual server IDs. Virtual server IDs are not technically equivalent to the membership IDs, because any trader may possess several virtual servers. However, we can define effective trader IDs, referred to as trading desks, by tracking all the virtual server IDs (see End Matter, the SM [9], and previous studies [27, 28, 29]). In this article, we refer to these trading desks as trader IDs for short.
@@
Figure 2: Stock-level price impact. (a) Average price-impact profile for four stocks (red: Toyota Motor Corp., blue: Nippon Telegraph and Telephone (NTT) Corp., green: SoftBank Group Corp., orange: Takeda Pharmaceutical Company Limited). Errorbars represent the standard errors of the means (SEMs). (b) Aggregate scaling plot, visualising the average scaling profiles across all stocks, confirms the square-root law for all stocks. (c) Histogram of the exponent $\delta$ across all stocks, exhibiting a sharp peak around $0.5$ with $\langle\delta\rangle=0.489$ with its SEM of $0.0015$ and $\sigma_{\delta}:=\sqrt{\langle(\delta-\langle\delta\rangle)^{2}\rangle}=0.071$. (d) Histogram of the prefactor $c$ across all stocks, with $\langle c\rangle=0.842$.
@@
We measured the transaction volume $Q$ and corresponding price impact $\Delta p$ for all liquid stocks on the TSE. First, we defined trader IDs for each stock and extracted the market-order volume sequences. Second, we assume that institutional investors typically split their large orders (called metaorders) into small orders (called child orders), as suggested by the literature [30, 28, 29, 31, 32]. Specifically, if two successive market orders have the same sign (plus for buy and minus for sell), we assume that both orders are child orders belonging to the same metaorder. The total volume $Q$ is then defined as the sum of all child-order volumes. Third, the price impact $\Delta p$ is defined as the price movement caused by such a metaorder. In this article, liquid stocks are defined as stocks where the total number of metaorders exceeds $10^{5}$.
@@
Figure 2(a) presents the typical price-impact profile $I(Q)$ for four liquid stocks. Additionally, we present Fig. 2(b) as the aggregate scaling plot—visualising the averaged impact profile across all stocks as one figure. The aggregate scaling plot was created by two steps (see End Matter and the SM [9]): (i) After non-dimensionalising $\mathcal{Q}:=Q/V_{\rm D}$ and $\mathcal{I}(\mathcal{Q}):=I(\mathcal{Q})/\sigma_{\rm D}$ for each stock, we applied fitting $\mathcal{I}(\mathcal{Q})=c\mathcal{Q}^{1/2}$ to estimate $c$. (ii) The average of $\mathcal{I}(\mathcal{Q})/c$ across all stocks was plotted, which collapsed onto a single master curve as theoretically expected. These two figures provide an overall picture of the SRL (1) and qualitatively support its validity across all stocks.
We next present the normalised histogram of $\delta$ across all stocks as Fig. 2(c). We estimated $\delta$ and $c$ using the power-law scaling (1) by applying the nonlinear relative least squares on binned averages of price impact (see SM [9] for a robustness check on the estimation method). Due to the finite-sample size, the $\delta$’s errorbars for individual stocks are roughly given by $\langle\!\langle\overline{\sigma_{\delta}}\rangle\!\rangle=0.063$, estimated by a plausible statistical model as discussed in the next paragraph. Figure 2(c) shows a very sharp peak around $1/2$ with the average and the standard deviation,
$\langle\delta\rangle=0.489\pm 0.0015,\>\>\>\sigma_{\delta}:=\sqrt{\langle(\delta-\langle\delta\rangle)^{2}\rangle}=0.071,$ (3)
@@
where $\langle A\rangle$ represents the cross-sectional average of AA on the TSE dataset across all stocks [^1] If we estimate $\delta$ by the nonlinear ordinary least squares on raw datapoints (without using binned averages), we obtain $\langle\delta_{\rm OLS}\rangle=0.474\pm 0.0012$. . This is the first main result of this article, suggesting that the exact SRL holds for more than two thousand data points. Also, Fig. 2(d) presents the histogram of the prefactor $c$. Furthermore, the crossover from the linear to square-root laws was observed consistently with Refs. [33, 4] (see SM [9]).
Errorbar estimation.
Statistically, estimating errorbars is necessary for $\delta$ due to finite sample size. If all the observations of $(\mathcal{Q},\mathcal{I})$ were independent and identically distributed (IID), simple nonlinear regression would suffice to estimate them. However, our dataset is a time series with serial correlation. For non-IID observations, estimating errorbars requires plausible statistical models to account for serial correlation (e.g., the autoregressive integrated moving average (ARIMA) model [34]). Thus, developing an appropriate statistical model for our dataset is crucial for obtaining reliable error estimates [^2] If we ignore the non-IID nature of the data and use standard error estimation packages (e.g., scipy.optimize.least_squares on Python [35]), the errorbar of $\delta$ is estimated to be $0.022$ via the nonlinear least relative squares on binned averages, or $0.014$ via the nonlinear ordinary least squares on raw datapoints. Both of these values are underestimates of the true errorbar in the presence of serial correlation. .
Figure 3: Errorbar estimation based on our statistical model. (a) $\delta$’s histgram across all stocks for one Monte Carlo iteration, consistent with the $\delta$’s histgram 2(c) for our TSE dataset. (b) Histgram of $\delta$’s errorbars $\langle\!\langle\sigma_{\delta}\rangle\!\rangle$ for individual stocks by the repeated Monte Carlo simulation. Each errorbar was estimated with $100$ iterations.
For this purpose, we develop a simple statistical model with the exact SRL under order splitting and estimate the corresponding errorbar (see End Matter and the SM [9]): (i) We randomly shuffled the signs of metaorders, while keeping the physical times of market-order submissions identical to those in our TSE dataset. (ii) The price impacts were assumed to exactly obey the SRL with $\delta=1/2$, with random contributions. The model parameters were based on our dataset. (iii) We repeatedly generated price time series for 100 iterations and measured $\delta$ for all stocks (see Fig. 3(a) for a $\delta$’s histogram for one numerical iteration, corresponding to Fig. 2(c) for our TSE dataset). (iv) The average and standard deviation of $\delta$ were evaluated across all Monte Carlo trials as $\langle\!\langle\overline{\delta}\rangle\!\rangle$ and $\langle\!\langle\overline{\sigma_{\delta}}\rangle\!\rangle$. For our numerical simulation, $\langle\!\langle A\rangle\!\rangle$ represents the Monte Carlo average of AA over $N_{\rm MC}=100$ iterations, whereas $\overline{A}$ represents the average across stocks for one Monte Carlo iteration. This model generates plausible price time-series, consistent with diffusive price. Indeed, Ref. [36] develops a nonlinear price-impact model under order splitting, reminiscent of this statistical model, to show its consistency with various stylized facts.
Since $\delta=1/2$ exactly in this statistical model, the dispersion $\langle\!\langle\overline{\sigma_{\delta}}\rangle\!\rangle$ represents the errorbar due to the finite sample size in the frequentist sense. Finally, we obtained
$\langle\!\langle\overline{\delta}\rangle\!\rangle=0.489\pm 0.0013,\>\>\>\langle\!\langle\overline{\sigma_{\delta}}\rangle\!\rangle=0.063.$ (4)
Hense, we estimate that our $\delta$’s errorbar is given by $\langle\!\langle\overline{\sigma_{\delta}}\rangle\!\rangle=0.063$ for the first main result in Fig. 2(c). Also, see Fig. 3(b) for the histogram of errorbars for individual stocks.
@@
Remarkably, the cross-sectional average $\langle\delta\rangle$ for the TSE dataset is equal to the Monte-Carlo average $\langle\!\langle\overline{\delta}\rangle\!\rangle$ for our statistical model within the errorbar, such that $|\langle\delta\rangle-\langle\!\langle\overline{\delta}\rangle\!\rangle|\ll\langle\!\langle\overline{\sigma_{\delta}}\rangle\!\rangle$. In addition, the cross-sectional dispersion $\sigma_{\delta}$ for the TSE dataset is almost equal to that $\langle\!\langle\overline{\sigma_{\delta}}\rangle\!\rangle$ for our statistical model, such that $\sigma_{\delta}\approx\langle\!\langle\overline{\sigma_{\delta}}\rangle\!\rangle$. These facts imply the self-consistency of our statistical analysis. Thus, we concluded that the standard deviation in estimating $\delta$ from our dataset is almost attributable to finite sample size, and the universality hypothesis holds for the SRL within statistical errors, at least for the TSE.
@@
Figure 4: Trader-level price impact. (a) Trader-level aggregate scaling plot confirms the square-root law for all traders. (b) Histogram of the exponent $\delta^{(i)}$ across all traders for $0\leq\delta^{(i)}\leq 1$ shows a sharp peak around $0.5$, with $\langle\delta^{(i)}\rangle=0.493$ with its SEM of $0.0050$ and $\sigma_{\delta^{(i)}}:=\sqrt{\langle(\delta^{(i)}-\langle\delta^{(i)}\rangle)^{2}\rangle}=0.177$. A few outliers were observed outside this region due to fitting errors. (c) Histogram of the prefactor $c^{(i)}$ across all traders, with $\langle c^{(i)}\rangle=1.501$.
While we presented evidence at the stock level, is the SRL universal even at the trader level? In other words, by writing the exponent for the $i$-th trader as $\delta^{(i)}$, is $\delta^{(i)}$ statistically indistinguishable from $1/2$ for all traders, despite the diversity in trading strategies among traders? To answer this question, we studied the price impact of all active traders—roughly defined as traders who submitted metaorders not less than $10^{4}$ times in this article (see End Matter and SM [9]). With this definition, we have 1,293 active traders in total. We present the aggregate scaling plot at the individual trader level as Fig. 4(a), qualitatively suggesting the validity of the SRL for all traders. As a second main result, we present the normalised trader-level histogram of $\delta^{(i)}$ as Fig. 4(b). The histogram has clear sharp peak with the average $\langle\delta^{(i)}\rangle$ across all traders on the TSE dataset and standard deviation $\sigma_{\delta^{(i)}}:=\sqrt{\langle(\delta^{(i)}-\langle\delta^{(i)}\rangle)^{2}\rangle}$ as
@@
The standard deviation $\sigma_{\delta^{(i)}}$ in our data was at the same level as that evaluated by our numerical model, $\langle\!\langle\overline{\delta^{(i)}}\rangle\!\rangle=0.521\pm 0.0048$ and $\langle\!\langle\overline{\sigma_{\delta^{(i)}}}\rangle\!\rangle=0.169$, where $\overline{A}$ represents the average across all traders in our numerical simulation. Since $|\langle\delta^{(i)}\rangle-\langle\!\langle\overline{\delta^{(i)}}\rangle\!\rangle|\ll\langle\!\langle\overline{\sigma_{\delta^{(i)}}}\rangle\!\rangle$ and $\sigma_{\delta^{(i)}}\approx\langle\!\langle\overline{\sigma_{\delta^{(i)}}}\rangle\!\rangle$, we concluded that the universality hypothesis of the SRL is maintained within statistical errors even at the trader level.